OpenStudy (berrymox):

Find the tangent line to the equation: x^4 = 4(x^2 y^2) at [1, sqrt(3/2) ] I got the derivative as (8y-4x^3)/8y Trying to simplify that seems to give me different slope, and the answer in the book is wrong. The book says the slop should be sqrt(3)

OpenStudy (berrymox):

\[y' = \frac{ 8x-4x ^{3} }{ 8y }\] \[f(x): x ^{4} = 4(x ^{2} y^{2})\]

zepdrix (zepdrix):

Hmm this problem is strange... Why not just divide both sides by x^2, \(\large\rm x^2=4y^2,\qquad\qquad x\ne0\) And differentiate from there. Thinking...

zepdrix (zepdrix):

Anyway, doing it the hard way :) \(\large\rm 4x^3=8xy^2+8x^2yy'\) Is this what you ended up with after differentiating implicitly?

OpenStudy (nincompoop):

can you show your work?

OpenStudy (robtobey2):

Refer to the Mathematica attachment.